Find the time periods of a simple pendulum of 1 metre length, placed on Earth and
th on Moon. The value of g on the surface of Moon is 1/6 of its value on Earth,
-2 where ge is 10 m s .
1
Expert's answer
2020-08-19T12:57:15-0400

The period of a simple pendulum on Earth is


"T_E=2\pi\sqrt{\frac{l}{g_E}}=2\pi\sqrt{\frac{1}{10}}=1.99\text{ s}."

The period of a simple pendulum on Moon is


"T_M=2\pi\sqrt{\frac{l}{(g_E/6)}}=2\pi\sqrt{\frac{1}{(10/6)}}=4.87\text{ s}."